solve:0.1p+2.1=4.6p-1.9+0.5p
Answer by jai_kos(139)

Given 0.1p + 2.1= 4.6p -1.9 + 0.5p
Group all the "p" terms to one side and the numbers on one side, we get..
0.1p - 4.6p - 0.5p = - 2.1 -1.9
0.1p -5.1p = - 4.0
- 5.0p = - 4. 0
p = 4 /5